\begin{thebibliography}{37}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{\citenamefont{Mohapatra and Valle}(1986)}]{Mohapatra:1986bd}
\bibinfo{author}{\bibfnamefont{R.~N.} \bibnamefont{Mohapatra}}
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.~W.~F.}
  \bibnamefont{Valle}}, \bibinfo{journal}{Phys. Rev. D}
  \textbf{\bibinfo{volume}{34}}, \bibinfo{pages}{1642} (\bibinfo{year}{1986}).

\bibitem[{\citenamefont{Wyler and Wolfenstein}(1983)}]{Wyler:1982dd}
\bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Wyler}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{L.}~\bibnamefont{Wolfenstein}},
  \bibinfo{journal}{Nucl. Phys. B} \textbf{\bibinfo{volume}{218}},
  \bibinfo{pages}{205} (\bibinfo{year}{1983}).

\bibitem[{\citenamefont{Fernandez-Martinez
  et~al.}(2016)\citenamefont{Fernandez-Martinez, Hernandez-Garcia, and
  Lopez-Pavon}}]{Fernandez-Martinez:2016lgt}
\bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Fernandez-Martinez}},
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Hernandez-Garcia}},
  \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Lopez-Pavon}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{08}}, \bibinfo{pages}{033}
  (\bibinfo{year}{2016}), \eprint{1605.08774}.

\bibitem[{\citenamefont{Agostinho et~al.}(2018)\citenamefont{Agostinho, Branco,
  Pereira, Rebelo, and Silva-Marcos}}]{Agostinho:2017wfs}
\bibinfo{author}{\bibfnamefont{N.~R.} \bibnamefont{Agostinho}},
  \bibinfo{author}{\bibfnamefont{G.~C.} \bibnamefont{Branco}},
  \bibinfo{author}{\bibfnamefont{P.~M.~F.} \bibnamefont{Pereira}},
  \bibinfo{author}{\bibfnamefont{M.~N.} \bibnamefont{Rebelo}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{J.~I.}
  \bibnamefont{Silva-Marcos}}, \bibinfo{journal}{Eur. Phys. J. C}
  \textbf{\bibinfo{volume}{78}}, \bibinfo{pages}{895} (\bibinfo{year}{2018}),
  \eprint{1711.06229}.

\bibitem[{\citenamefont{Das et~al.}(2017)\citenamefont{Das, Nomura, Okada, and
  Roy}}]{Das:2017ski}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Das}},
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Roy}},
  \bibinfo{journal}{Phys. Rev. D} \textbf{\bibinfo{volume}{96}},
  \bibinfo{pages}{075001} (\bibinfo{year}{2017}), \eprint{1704.02078}.

\bibitem[{\citenamefont{Nomura and Okada}(2019{\natexlab{a}})}]{Nomura:2018cfu}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}},
  \bibinfo{journal}{Phys. Rev. D} \textbf{\bibinfo{volume}{99}},
  \bibinfo{pages}{055027} (\bibinfo{year}{2019}{\natexlab{a}}),
  \eprint{1807.04555}.

\bibitem[{\citenamefont{Dey et~al.}(2019)\citenamefont{Dey, Nomura, and
  Okada}}]{Dey:2019cts}
\bibinfo{author}{\bibfnamefont{U.~K.} \bibnamefont{Dey}},
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}},
  \bibinfo{journal}{Phys. Rev. D} \textbf{\bibinfo{volume}{100}},
  \bibinfo{pages}{075013} (\bibinfo{year}{2019}), \eprint{1902.06205}.

\bibitem[{\citenamefont{Nomura and Okada}(2020)}]{Nomura:2020cog}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}
  (\bibinfo{year}{2020}), \eprint{2007.15459}.

\bibitem[{\citenamefont{Nomura and Okada}(2019{\natexlab{b}})}]{Nomura:2018ktz}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}},
  \bibinfo{journal}{Phys. Lett. B} \textbf{\bibinfo{volume}{792}},
  \bibinfo{pages}{424} (\bibinfo{year}{2019}{\natexlab{b}}),
  \eprint{1809.06039}.

\bibitem[{\citenamefont{Nomura and Okada}(2024)}]{Nomura:2024jxc}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}
  (\bibinfo{year}{2024}), \eprint{2403.14193}.

\bibitem[{\citenamefont{Nomura et~al.}(2022)\citenamefont{Nomura, Okada, and
  Sanyal}}]{Nomura:2021adf}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Sanyal}},
  \bibinfo{journal}{Eur. Phys. J. C} \textbf{\bibinfo{volume}{82}},
  \bibinfo{pages}{697} (\bibinfo{year}{2022}), \eprint{2103.09494}.

\bibitem[{\citenamefont{Okada and Toma}(2012)}]{Okada:2012np}
\bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Toma}},
  \bibinfo{journal}{Phys. Rev. D} \textbf{\bibinfo{volume}{86}},
  \bibinfo{pages}{033011} (\bibinfo{year}{2012}), \eprint{1207.0864}.

\bibitem[{\citenamefont{Kobayashi
  et~al.}(2025{\natexlab{a}})\citenamefont{Kobayashi, Okada, and
  Otsuka}}]{Kobayashi:2025cwx}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}
  (\bibinfo{year}{2025}{\natexlab{a}}), \eprint{2505.14878}.

\bibitem[{\citenamefont{Okada and Shigekami}(2025)}]{Okada:2025kfm}
\bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Shigekami}}
  (\bibinfo{year}{2025}), \eprint{2507.16198}.

\bibitem[{\citenamefont{Choi et~al.}(2022)\citenamefont{Choi, Lam, and
  Shao}}]{Choi:2022jqy}
\bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Choi}},
  \bibinfo{author}{\bibfnamefont{H.~T.} \bibnamefont{Lam}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.-H.} \bibnamefont{Shao}},
  \bibinfo{journal}{Phys. Rev. Lett.} \textbf{\bibinfo{volume}{129}},
  \bibinfo{pages}{161601} (\bibinfo{year}{2022}), \eprint{2205.05086}.

\bibitem[{\citenamefont{Cordova
  et~al.}(2024{\natexlab{a}})\citenamefont{Cordova, Hong, Koren, and
  Ohmori}}]{Cordova:2022fhg}
\bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Cordova}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Hong}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Koren}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{K.}~\bibnamefont{Ohmori}},
  \bibinfo{journal}{Phys. Rev. X} \textbf{\bibinfo{volume}{14}},
  \bibinfo{pages}{031033} (\bibinfo{year}{2024}{\natexlab{a}}),
  \eprint{2211.07639}.

\bibitem[{\citenamefont{Cordova and Ohmori}(2023)}]{Cordova:2022ieu}
\bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Cordova}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{K.}~\bibnamefont{Ohmori}},
  \bibinfo{journal}{Phys. Rev. X} \textbf{\bibinfo{volume}{13}},
  \bibinfo{pages}{011034} (\bibinfo{year}{2023}), \eprint{2205.06243}.

\bibitem[{\citenamefont{Cordova
  et~al.}(2024{\natexlab{b}})\citenamefont{Cordova, Hong, and
  Koren}}]{Cordova:2024ypu}
\bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Cordova}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Hong}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Koren}}
  (\bibinfo{year}{2024}{\natexlab{b}}), \eprint{2402.12453}.

\bibitem[{\citenamefont{Kobayashi et~al.}(2024)\citenamefont{Kobayashi, Otsuka,
  and Tanimoto}}]{Kobayashi:2024cvp}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Tanimoto}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{12}}, \bibinfo{pages}{117}
  (\bibinfo{year}{2024}), \eprint{2409.05270}.

\bibitem[{\citenamefont{Kobayashi and Otsuka}(2024)}]{Kobayashi:2024yqq}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{11}}, \bibinfo{pages}{120}
  (\bibinfo{year}{2024}), \eprint{2408.13984}.

\bibitem[{\citenamefont{Kobayashi
  et~al.}(2025{\natexlab{b}})\citenamefont{Kobayashi, Nishioka, Otsuka, and
  Tanimoto}}]{Kobayashi:2025znw}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Nishioka}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Tanimoto}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{05}}, \bibinfo{pages}{177}
  (\bibinfo{year}{2025}{\natexlab{b}}), \eprint{2503.09966}.

\bibitem[{\citenamefont{Suzuki and Xu}(2025)}]{Suzuki:2025oov}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Suzuki}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{L.-X.} \bibnamefont{Xu}}
  (\bibinfo{year}{2025}), \eprint{2503.19964}.

\bibitem[{\citenamefont{Liang and Yanagida}(2025)}]{Liang:2025dkm}
\bibinfo{author}{\bibfnamefont{Q.}~\bibnamefont{Liang}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{T.~T.} \bibnamefont{Yanagida}}
  (\bibinfo{year}{2025}), \eprint{2505.05142}.

\bibitem[{\citenamefont{Kobayashi
  et~al.}(2025{\natexlab{c}})\citenamefont{Kobayashi, Otsuka, Tanimoto, and
  Uchida}}]{Kobayashi:2025ldi}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}},
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Tanimoto}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Uchida}}
  (\bibinfo{year}{2025}{\natexlab{c}}), \eprint{2505.07262}.

\bibitem[{\citenamefont{Kobayashi
  et~al.}(2025{\natexlab{d}})\citenamefont{Kobayashi, Mita, Otsuka, and
  Sakuma}}]{Kobayashi:2025lar}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Mita}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Sakuma}}
  (\bibinfo{year}{2025}{\natexlab{d}}), \eprint{2506.10241}.

\bibitem[{\citenamefont{Nomura and Okada}(2025)}]{Nomura:2025sod}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}
  (\bibinfo{year}{2025}), \eprint{2506.16706}.

\bibitem[{\citenamefont{Dong et~al.}(2025)\citenamefont{Dong, Jeric, Kobayashi,
  Nishida, and Otsuka}}]{Dong:2025jra}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Dong}},
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Jeric}},
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Nishida}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}
  (\bibinfo{year}{2025}), \eprint{2507.02375}.

\bibitem[{\citenamefont{Nomura and Popov}(2025)}]{Nomura:2025yoa}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nomura}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{O.}~\bibnamefont{Popov}}
  (\bibinfo{year}{2025}), \eprint{2507.10299}.

\bibitem[{\citenamefont{Chen et~al.}(2025)\citenamefont{Chen, Geng, Okada, and
  Wu}}]{Chen:2025awz}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Chen}},
  \bibinfo{author}{\bibfnamefont{C.-Q.} \bibnamefont{Geng}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Okada}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.-J.} \bibnamefont{Wu}}
  (\bibinfo{year}{2025}), \eprint{2507.11951}.

\bibitem[{\citenamefont{Kobayashi
  et~al.}(2025{\natexlab{e}})\citenamefont{Kobayashi, Otsuka, and
  Yanagida}}]{Kobayashi:2025thd}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Kobayashi}},
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Otsuka}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{T.~T.} \bibnamefont{Yanagida}}
  (\bibinfo{year}{2025}{\natexlab{e}}), \eprint{2508.12287}.

\bibitem[{\citenamefont{Suzuki et~al.}(2025)\citenamefont{Suzuki, Xu, and
  Zhang}}]{Suzuki:2025bxg}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Suzuki}},
  \bibinfo{author}{\bibfnamefont{L.-X.} \bibnamefont{Xu}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.~Y.} \bibnamefont{Zhang}}
  (\bibinfo{year}{2025}), \eprint{2508.14970}.

\bibitem[{\citenamefont{Chang et~al.}(2019)\citenamefont{Chang, Lin, Shao,
  Wang, and Yin}}]{Chang:2018iay}
\bibinfo{author}{\bibfnamefont{C.-M.} \bibnamefont{Chang}},
  \bibinfo{author}{\bibfnamefont{Y.-H.} \bibnamefont{Lin}},
  \bibinfo{author}{\bibfnamefont{S.-H.} \bibnamefont{Shao}},
  \bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Wang}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{X.}~\bibnamefont{Yin}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{01}}, \bibinfo{pages}{026}
  (\bibinfo{year}{2019}), \eprint{1802.04445}.

\bibitem[{\citenamefont{Ma}(2006)}]{Ma:2006km}
\bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Ma}}, \bibinfo{journal}{Phys.
  Rev. D} \textbf{\bibinfo{volume}{73}}, \bibinfo{pages}{077301}
  (\bibinfo{year}{2006}), \eprint{hep-ph/0601225}.

\bibitem[{\citenamefont{Kanemura et~al.}(2010)\citenamefont{Kanemura,
  Matsumoto, Nabeshima, and Okada}}]{Kanemura:2010sh}
\bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Kanemura}},
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Matsumoto}},
  \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Nabeshima}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Okada}},
  \bibinfo{journal}{Phys. Rev. D} \textbf{\bibinfo{volume}{82}},
  \bibinfo{pages}{055026} (\bibinfo{year}{2010}), \eprint{1005.5651}.

\bibitem[{\citenamefont{Aghanim et~al.}(2020)}]{Planck:2018vyg}
\bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Aghanim}} \bibnamefont{et~al.}
  (\bibinfo{collaboration}{Planck}), \bibinfo{journal}{Astron. Astrophys.}
  \textbf{\bibinfo{volume}{641}}, \bibinfo{pages}{A6} (\bibinfo{year}{2020}),
  \bibinfo{note}{[Erratum: Astron.Astrophys. 652, C4 (2021)]},
  \eprint{1807.06209}.

\bibitem[{\citenamefont{Abe et~al.}(2024)}]{KamLAND-Zen:2024eml}
\bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Abe}} \bibnamefont{et~al.}
  (\bibinfo{collaboration}{KamLAND-Zen}) (\bibinfo{year}{2024}),
  \eprint{2406.11438}.

\bibitem[{\citenamefont{Esteban et~al.}(2024)\citenamefont{Esteban,
  Gonzalez-Garcia, Maltoni, Martinez-Soler, Pinheiro, and
  Schwetz}}]{Esteban:2024eli}
\bibinfo{author}{\bibfnamefont{I.}~\bibnamefont{Esteban}},
  \bibinfo{author}{\bibfnamefont{M.~C.} \bibnamefont{Gonzalez-Garcia}},
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Maltoni}},
  \bibinfo{author}{\bibfnamefont{I.}~\bibnamefont{Martinez-Soler}},
  \bibinfo{author}{\bibfnamefont{J.~a.~P.} \bibnamefont{Pinheiro}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Schwetz}},
  \bibinfo{journal}{JHEP} \textbf{\bibinfo{volume}{12}}, \bibinfo{pages}{216}
  (\bibinfo{year}{2024}), \eprint{2410.05380}.

\end{thebibliography}

\documentclass[prd,preprint,superscriptaddress,amsmath,amssymb,nofootinbib]{revtex4}
\usepackage{graphicx}
\usepackage{dcolumn}
\usepackage{bm}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{epsfig}    
\usepackage{color}
\usepackage{slashed}
\usepackage{hhline}
\usepackage{bbm}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}

\begin{document}

\title{A natural realization of inverse seesaw model in a non-invertible selection rule}

\author{Shilpa Jangid }
\email{shilpajangid123@gmail.com}
\affiliation{Shiv Nadar IoE Deemed to be University, Gautam Buddha Nagar, Uttar Pradesh, 201314, India}

\author{ Hiroshi Okada}
\email{hiroshi3okada@htu.edu.cn}
\affiliation{Department of Physics, Henan Normal University, Xinxiang 453007, China}

\date{\today}

\begin{abstract}
We propose natural hierarchies among neutral fermions in a framework of inverse seesaw, imposing a $Z_3$ Tambara-Yamagami fusion rule which is applied to our phenomenology. Under the symmetry, the Majorana mass terms for $N_R$ and $N_L$ are forbidden at tree level.  However they are generated at one-loop level where the symmetry is dynamically broken.
In order to realize such loop corrections, we introduce neutral boson and fermion either of which can be an appropriate dark matter candidate.
Finally, we show the best fit value of the neutrino sector for normal and inverted hierarchies referring recent experimental results.

 \end{abstract}
\maketitle

\section{Introduction}
Inverse seesaw is a verifiable scenario of generating tiny neutrino masses within TeV scale~\cite{Mohapatra:1986bd, Wyler:1982dd}.
The model requests two types of neutral fermions $N_R$ and $N_L$, where $N_R$ has right-handed chirality and $N_L$ has left-handed one.
Then, we generally have the following mass terms in collaboration with the standard model (SM) particles ; 
\begin{align}
m_{D_R} {\overline N_R} \nu_L +m_{D_L} {\overline N_L^C} \nu_L +
M_D {\overline N_R}N_L + m_R {\overline N_R} N_R^C +m_L {\overline N_L^C} N_L +{\rm c.c.}\ , \label{eq:iss}
\end{align}
where $m_{D_R}\equiv y_{D_R} v_H/\sqrt2$ and $m_{D_L}\equiv y_{D_L} v_H/\sqrt2$ are respectively obtained after spontaneous symmetry breaking $y_{D_R}{\overline N_R} \tilde H L_L$ $y_{D_L}{\overline N_L^C} \tilde H L_L$.
$H$ is the SM Higgs boson whose vacuum expectation value (VEV) is denoted by $\langle H\rangle\equiv [0,v_H/\sqrt2]^T$, and 
$\tilde H\equiv i\sigma_2 H^*$, $\sigma_2$ being the second Pauli matrix.
Then, the mass matrix for the neutral fermions in basis of $[\nu_L,N^C_R,N_L]^T$ is given by
\begin{align}
\begin{pmatrix}
 0 & m_{D_R}^T & m_{D_L}^T  \\
 m_{D_R} & m_R & M_D \\
 m_{D_L} & M_D^T & m_L 
  \end{pmatrix}.
\end{align}
When we impose the following mass hierarchies among them~\footnote{There are several ways of the hierarchies to realize the inverse seesaw. }
\begin{align}
 m_{D_L} \ll
   m_R  , m_L \ll m_{D_R} \lesssim  M_D, \label{eq:mass-order}
\end{align}
the active neutrino mass matrix is found as
\begin{align}
m_\nu\simeq m^T_{D_R} (M_D^T)^{-1} m_L M_D^{-1} m_{D_R}, \label{eq:iss}
\end{align}
which is called Inverse Seesaw (ISS).
%
Since the active neutrinos mix with the heavier neutral fermion sector, we need to consider constraints from non-unitarity.
The constraints can be characterized by the hermitian matrix $(F\equiv)  -M_D^{-1} m_{D_R}$, and
the non-unitarity matrix, which is denoted by $U'$, is given by the following form 
\begin{align}
U'\equiv \left(1-\frac12 FF^\dag\right) U.
\end{align}
$U$ is a unitary matrix to diagonalize the active neutrino mass matrix $m_\nu$. Moreover, $U$ can be testable one that involves several observables such as mixing angles and Dirac CP phase under the diagonal mass matrix for charged-leptons. In fact, we work on this setup in our main discussion below. 
%
Considering several experimental results such as the effective Weinberg angle, SM $W$ boson mass, several ratios of $Z$ boson fermionic decays, invisible decay of $Z$, electroweak universality, measurements of the quark mixing matrix, and lepton flavor violations~\cite{Fernandez-Martinez:2016lgt},
the most stringent constraint comes from 1-2 component of $|FF^\dag|$~\cite{Agostinho:2017wfs, Das:2017ski} and  is found as
\begin{align}
|F| \lesssim 4.90\times 10^{-3}. \label{eq:f}
\end{align}
It suggests that $m_{D_R}\ll M_D$ is experimentally assured in Eq.~(\ref{eq:mass-order}).
In a theoretical viewpoint, $M_D$ can be expected to originate from new scale. When we suppose our new physics is within TeV scale,
$M_D$ would be ${\cal O}(1)$ TeV. On the other hand, $m_{D_R}(\sim v_H y_{D_R})$ is generated after the spontaneous electroweak symmetry breaking whose scale is $v_H\sim {\cal O}$(100) GeV.
Thus, if the Yukawa coupling is $y_{D_R}\sim {\cal O}(0.01)$, the required hierarchy in Eq.~(\ref{eq:f}) can be achieved.
However we need to introduce an additional symmetry or new particles theoretically to understand the following hierarchies
\begin{align} m_{D_L} \ll  m_R  , m_L \ll m_{D_R}.  \label{eq:mass-order-iss} \end{align}
%
For examples, refs~\cite{ Nomura:2018cfu,  Dey:2019cts, Nomura:2020cog} show that $m_{R,L}$ are generated via VEV of multiplet boson  that is restricted by the $\rho$ parameter and its VEV has to be less than ${\cal O}(1)$ GeV while $m_{D_L}$ is forbidden by an additional symmetry. 
Thus $m_{R,L}\ll m_{D_{R}}$ is theoretically explained.
On the other hand, refs.~\cite{Nomura:2018ktz, Nomura:2024jxc, Nomura:2021adf, Okada:2012np} demonstrate that $m_{R,L}$ are induced via one-loop level while forbidding $m_{D_L}$ via an additional symmetry. Hence, $m_{R,L}\ll m_{D_{R}}$ is explained well.

In this paper, we propose a reasonable explanation among these hierarchies in Eq.~(\ref{eq:mass-order-iss}) via corrections,
introducing a rather new symmetry; $Z_3$ Tambara-Yamagami fusion rule ($Z_3$-TY). This is recently applied to phenomenologies~\cite{Kobayashi:2025cwx}.
\footnote{One find several literatures to apply these kinds of symmetries to phenomenologies~\cite{Okada:2025kfm, Choi:2022jqy,Cordova:2022fhg,Cordova:2022ieu,Cordova:2024ypu,Kobayashi:2024cvp,Kobayashi:2024yqq,Kobayashi:2025znw,Suzuki:2025oov,Liang:2025dkm,Kobayashi:2025ldi,Kobayashi:2025lar,Kobayashi:2025cwx,Nomura:2025sod,Dong:2025jra,Nomura:2025yoa,Chen:2025awz, Kobayashi:2025thd, Suzuki:2025bxg}.}
$Z_3$-TY selection rules are simply given by~\cite{Dong:2025jra, Chang:2018iay}
\begin{align}
a\otimes a =a^*,\quad a^*\otimes a^* =a,\quad a\otimes a^*={\mathbbm I},\quad a\otimes  n=a^*\otimes n=n,\quad n\otimes  n={\mathbbm I}\oplus a\oplus a^*,
\end{align}
where $a^{(*)}$ and $n$ are commutable each other.
This symmetry possesses several good nature to have model buildings; especially, the dynamical symmetry breaking mechanism~\cite{Kobayashi:2025cwx}.
This mechanism provides us radiative couplings or masses at loop level while tree level ones are forbidden by this symmetry.
~\footnote{In fact, one can straightforwardly derive Ma model~\cite{Ma:2006km} via this mechanism. It is currently known that Ma model of this mechanism can be constructed by 
only two selection rules; $Z_3$-TY and $Z_3$ gauging of $\mathbbm{Z}_3\times \mathbbm{Z'}_3$~\cite{Dong:2025jra}.} 
Applying this mechanism to the ISS, the terms of $m_{R,L}$ are forbidden at tree level but generated at one-loop level.
While $m_{D_L}$ is totally forbidden by this symmetry. In addition, dark matter (DM) candidate can be involved in this model and the DM contributes to generate the $m_{R,L}$
as can be seen in the next section.

This paper is organized as follows. In Section~II, we present the model with mass spectrum, relevant interactions and formulas for phenomenology of our interest.
In Section III, we perform numerical analysis to search for allowed parameter region and predicted flavor observables. 
Finally we devote Section~IV to the summary and conclusion.

\section{Model setup}
\label{sec:II}

\begin{table}[t!]
\begin{tabular}{|c||c|c|c|c|c||c|c|}\hline\hline  
& ~$L_L$~& ~$\ell_R$~ & ~$N_L$~ & ~${N_R}$~ & ~${X_R}$~ & ~$H$~  & ~{$S$}~    \\\hline\hline 
$SU(2)_L$   & $\bm{2}$  & $\bm{1}$  & $\bm{1}$ & $\bm{1}$ & $\bm{1}$   & $\bm{2}$      & $\bm{1}$  \\\hline 
$U(1)_Y$    & $-\frac12$  & $-1$  & $0$  & $0$    & $0$ & $\frac12$      & $0$    \\\hline
{$Z_3$-TY}   & ${a}$  & $ a $& $ a$ & $a$ & $n $  & $\mathbbm{I}$ & $n$ \\\hline 
 \end{tabular}
\caption{Charge assignments of the fields under the $SU(2)_L \times U(1)_Y$ gauge symmetry and the IFR.
}\label{tab:1}
\end{table}

Here, we show our model.
In addition to the ISS particles in Eq.~(\ref{eq:iss}), we introduce three right-handed neutral fermions $X_R$ and an inert singlet boson $S\equiv [S_R + S_I]/\sqrt2$. These two particles are assigned by $n$ while the other particles except $H$ are assigned by $a$.
The SM Higgs is neutral under {$Z_3$-TY}. All the particle contents and their assignments are listed in Table~\ref{tab:1}.
Under this symmetry, $m_{R,L}$ are forbidden at tree level since ${\mathbbm I}\notin a^2$ and ${\mathbbm I}\notin  (a^*)^2$.
Below we show forbidden renormalizable terms due to  {$Z_3$-TY};
\begin{align}
& {\overline N^C_L} N_L,\quad {\overline N^C_R} N_R,\quad {\overline N^C_L} N_L S,\quad {\overline N^C_R} N_R S, \quad 
{\overline N_L} X_R,\quad {\overline L_L} \tilde H N^C_L,\quad \\
& {\overline N^C_R} X_R,\quad {\overline L_L} \tilde H X_R , \quad S^3.
\end{align}
{\it Note here that vanishing the term of $S^3$ is crucial for model building. Otherwise, there is logarithmic divergence from the $S$ decay into two Higgses at one-loop level that spoils our model. Thus, $Z_3$ gauging ${\mathbbm Z}_3\times {\mathbbm Z'}_3$ does not work well by itself because of non-vanishing term of $S^3$.} 
But all the other terms in Eq.~(\ref{eq:iss}) are allowed.
Then, $N_{L,R}$ can have the following interactions which are invariant under {$Z_3$-TY} symmetry %
\begin{align}
f_{L_{\alpha a}} \overline{X_{R_\alpha}} N_{L_a} S 
+
f_{R_{a\alpha}} \overline{N_{R_a}} X^C_{R_\alpha} S
+
M_{X_\alpha}  \overline{X_{R_\alpha}} X_{R_\alpha}
-\mu^2_S S^2
+{\rm h.c.}. \label{eq:m_RL}
 \end{align}
 The above terms lead us to generate $m_{R,L}$ at one-loop level where {$Z_3$-TY} is dynamically broken.
 As a result, we find the following mass terms:
 \begin{align}
 m_{L_{ab}}&=
\frac1{2(4\pi)^2}
\sum_{\alpha=1}^3  
{f^T_{L_{a\alpha}} M_{X_\alpha} f_{L_{\alpha b}} }
\left[
\frac{m^2_{R}}{m^2_R -M_{X_\alpha}}\ln\left(\frac{m^2_{R}}{M_{X_\alpha}^2}\right)
-
\frac{m^2_{I}}{m^2_I - M_{X_\alpha}}\ln\left(\frac{m^2_{I}}{M_{X_\alpha}^2}\right)
\right], \\
 m_{R_{ab}}&=
\frac1{2(4\pi)^2}
\sum_{\alpha=1}^3  
{f_{R_{a\alpha}} M_{X_\alpha} f^T_{R_{\alpha b}} }
\left[
\frac{m^2_{R}}{m^2_R -M_{X_\alpha}}\ln\left(\frac{m^2_{R}}{M_{X_\alpha}^2}\right)
-
\frac{m^2_{I}}{m^2_I - M_{X_\alpha}}\ln\left(\frac{m^2_{I}}{M_{X_\alpha}^2}\right)
\right], \label{eq:m_lR-form}
\end{align}
where $m_R$ and $m_I$ are respectively the mass eigenvalues of $S_R$ and $S_I$.
Here, we redefine to be $m^2_0\equiv (m_R^2+m_I^2)/2$ and $\delta m^2\equiv (m_R^2 - m_I^2)/2$ where we suppose $\delta m^2\ll M_{X_\alpha}^2$. Furthermore, $r_\alpha \equiv m^2_0/M_{X_\alpha}^2$.
Then we rewrite them as follows:
 \begin{align}
 m_{L_{ab}}&\simeq-
\frac{\delta m}{(4\pi)^2}
\sum_{\alpha=1}^3  
 \frac{\delta m}{M_{X_\alpha}}
\frac{f^T_{L_{a\alpha}} f_{L_{\alpha b}}}{(1-r_\alpha)^2} 
\left(
1-r_\alpha+\ln[r_\alpha]
\right), \\
 m_{R_{ab}}&\simeq-
\frac{\delta m}{(4\pi)^2}
\sum_{\alpha=1}^3  
 \frac{\delta m}{M_{X_\alpha}}
\frac{f_{R_{a\alpha}} f^T_{R_{\alpha b}}}{(1-r_\alpha)^2} 
\left(
1-r_\alpha+\ln[r_\alpha]
\right). \label{eq:m_lR-approx}
\end{align}
Here, we briefly comment on our DM candidate.
The DM stability is assured by $n$ under the {$Z_3$-TY}, therefore, the lightest fermionic DM of $X_R$ or bosonic one, $S$.
However, since the fermionic DM does not have any interactions to explain the correct relic density, $S$ is the only channel to be the main source of the DM.
The detailed analysis has already been discussed in ref.~\cite{Kanemura:2010sh},
in which the allowed region to satisfy the observed relic density and bound on direct detection searches is at nearby half of the SM Higgs mass $m_H\approx126$ GeV or the DM mass is greater than ${\cal O}(1)$ TeV. 

The active neutrino mass matrix is given in Eq~(\ref{eq:iss}), where we rewrite each of the mass matrix as follows:
\begin{align}
M_D &=M_3
\left[\begin{array}{ccc}
\tilde M_1 & 0& 0\\
0 & \tilde M_2 & 0 \\
0 & 0 & 1 \\
\end{array}\right]\equiv M_3 \tilde M_D,\\
m_{D_R} &=\frac{v_H y_{33}}{\sqrt2}
\left[\begin{array}{ccc}
 y_{11} & y_{12} & y_{13}\\
y_{21} & y_{22} & y_{23}\\
y_{31} & y_{32} & 1\\
\end{array}\right]
\equiv \frac{v_H y_{33}}{\sqrt2} \tilde m_{D_R},\\
m_L &\equiv -\frac{\delta m}{(4\pi)^2}\tilde m_L.
\end{align}
 Then, $m_\nu$ is rewrite as 
\begin{align}
m_\nu
&\simeq -\frac{y^2_{33}}{2(4\pi)^2}\frac{v^2_H }{M_3^2}  \delta m 
\left[\tilde m^T_{D_R} (\tilde M_D)^{-1} \tilde m_L \tilde M_D^{-1} \tilde m_{D_R}\right]
\equiv \kappa_\nu\tilde m_\nu. \label{eq:iss_ours}
\end{align}
$m_\nu$ is diagonalized by a unitary matrix $U$, that is,  $D_\nu=U^T m_\nu U\equiv \kappa_\nu U^T \tilde m_\nu U$.
Furthermore $\tilde D_\nu \equiv{\rm diag}[\tilde D_1, \tilde D_2,\tilde D_3] \equiv D_\nu/\kappa_\nu$.
{\it $U$ is observed matrix where charged-lepton mass matrix can be diagonal without loss of generality.}
Then, two  squared mass differences $\Delta m^2_{atm}$ and $\Delta m^2_{sol}$ can be written in terms of our parameters as follows:
\begin{align}
\kappa^2_\nu=\frac{\Delta m^2_{sol}}{\tilde D_2^2-\tilde D_1^2}, \label{eq:kappa}
\end{align}
and
\begin{align}
&{\rm NH}: \Delta m^2_{atm} = \Delta m^2_{sol}\left( \frac{\tilde D_3^2-\tilde D_1^2}{\tilde D_2^2-\tilde D_1^2}\right),\\
&{\rm IH}: \Delta m^2_{atm} = \Delta m^2_{sol}\left( \frac{\tilde D_2^2-\tilde D_3^2}{\tilde D_2^2-\tilde D_1^2}\right).
\end{align}
Sum of neutrino masses; $\sum D_\nu=\kappa_\nu(\tilde D_1  +\tilde  D_2+\tilde  D_3)$, should be less 120 meV~\cite{Planck:2018vyg} coming from the minimal standard cosmological model with CMB data.
%
The effective mass for neutrinoless double beta decay $m_{ee}$ is defined as follows
\begin{align}
 m_{ee}= \kappa_\nu \left| \tilde D_{1} c^2_{12} c^2_{13}+ \tilde D_{2} s^2_{12} c^2_{13}e^{i\alpha}+ \tilde D_{3} s^2_{13}e^{i(\beta-2\delta_{CP})} \right|,
\end{align}
where $s_{12,23,13} (c_{12,23,13})$, which are short-hand notations $\sin\theta_{12,23,13} (\cos\theta_{12,23,13})$, are neutrino mixing of $U$, $\delta_{CP}$ is Dirac phase, $P\equiv {\rm diag}[1,e^{i\alpha/2},e^{i\beta/2}]$ are Majorana phases, $P$ being the Majorana phase matrix. 
The upper bound on $m_{ee}$ is given by the current KamLAND-Zen data measured in the future~\cite{KamLAND-Zen:2024eml}
that tells us $ m_{ee} <(36-156)$ meV at 90 \% confidence level (CL).

 \section{Numerical results}
 \label{sect.3}
 Here, we perform our numerical analysis and show benchmark points for NH and IH with the minimum chi square.
 Before that, we simplify $m_{D_R}$ to be the diagonal mass matrix; we set $y_{21}=y_{31}=y_{32}=y_{12}=y_{13}=y_{23}\sim 0$, and $\tilde m_{D_R}\sim {\rm diag}[y_{11},y_{22},1]$. $f_L$ is three by three complex dimensionless parameters, therefore we suppose that observed mixing angles and phases originates from $f_L$.
Since we simplify $m_{D_R}$, Max[$|\tilde M_D^{-1} \tilde m_{D_R}|$]$=1$. 
That is, the constraint of $|F|$ is found  as follows:
\begin{align}
& |F|=\left(\frac{v_H y_{33}}{\sqrt2 M_3}\right) |\tilde M_D^{-1} \tilde m_{D_R}| \lesssim 4.9\times 10^{-3},\\
&q_{min}\equiv \left(\frac{v_H y_{33}}{\sqrt2 M_3}\right)_{min}\lesssim \frac{4.9\times 10^{-3}}{ |\tilde M_D^{-1} \tilde m_{D_R}|_{max}}
=  4.9\times 10^{-3}.\label{eq:q}
\end{align}
On the other hand, since $\kappa_\nu$ is experimentally determined in Eq.~(\ref{eq:kappa}),
$q_{min}$ in Eq.~(\ref{eq:q}) is rewritten in terms of $\kappa_\nu$ and $\delta m$ and we find the following condition
\begin{align}
(4\pi) \sqrt{\frac{\kappa_\nu}{\delta m}} \lesssim  4.9\times 10^{-3}.
\end{align}
In addition to the above constraints, we adopt Nufit 6.0~\cite{Esteban:2024eli} as experimental observables and remind it $m_0\simeq m_S=m_h/2$.
We estimate $\chi^2$ by applying the following formula
\begin{align}
\chi^2 = \sum_{i}  \left( \frac{O_i^{\rm obs} - O_i^{\rm th}}{\delta O_i^{\rm exp}} \right)^2, \label{eq:chi-square}
\end{align}
where $O_i^{\rm obs (th)}$ is observed (theoretically) obtained value of corresponding observables and $\delta O_i^{\rm exp}$ indicates the experimental error at $1\sigma$. Here, we adopt five reliable observables; $s_{12},s_{23},s_{13},\Delta m^2_{sol},\ \Delta m^2_{atm}$, assuming all are Gaussian.
\\
In case of {\it NH} case, we obtain the best fit value of $\chi^2$ is 0.0593 and show our input values and observables in Table~\ref{BP-nh}.

 \begin{table}[h!]
 		\centering
 	\begin{tabular}{|c|c|} \hline 
 		\rule[14pt]{0pt}{0pt}	
 		$\delta m$& $7.66$ GeV  \\ 
 		\rule[14pt]{0pt}{0pt}
		$[M_{X_1},M_{X_2},M_{X_3}]$& $[862,\ 2.86\times10^4,\ 5.63\times10^4]$ GeV  \\ 
 		\rule[14pt]{0pt}{0pt}
 		$[\tilde M_{1},\tilde M_{2}]$ &$[0.0168,\ 0.143]$ \\
 		\rule[14pt]{0pt}{0pt}
 		$[y_{11},\ y_{22}]$ &$[0.000848,\ 0.100]$ \\
 		\rule[14pt]{0pt}{0pt}
 		$ f_L \times10^2$ & $
		\left[\begin{array}{ccc}
 -6.83 + 0.621 i & 4.29 + 2.88 i & 1.81 + 1.99 i \\
13.8 + 23.9 i & 1.32 - 3.30 i & 0.0398 + 0.00386 i \\
4.92 - 82.8 i & 3.14 - 15.3 i & 0.0310 - 0.0355 i \\
\end{array}\right]$  \\
 		\rule[14pt]{0pt}{0pt}
         	$|F|_{max}$ & $0.00382$   \\
 		\rule[14pt]{0pt}{0pt}
 	$\kappa_\nu$ & $ 7.08\times 10^{-7}$  \\
 		\rule[14pt]{0pt}{0pt}
	$[\sin\theta_{12},\ \sin\theta_{23},\ \sin\theta_{13}]$ & $ [0.550, 0.758, 0.149]$  \\
 		\rule[14pt]{0pt}{0pt}
 	$[\Delta m^2_{sol},\ \Delta m^2_{atm}] $ & $[7.39\times 10^{-5},\ 2.51\times 10^{-3}]{\rm eV^2}$  \\
 		\rule[14pt]{0pt}{0pt}
 		$[\delta_{CP},\alpha,\beta]$ & $[141^\circ, 343 ^\circ, 109 ^\circ]$ \\
 		\rule[14pt]{0pt}{0pt}
	$[m_{ee},\ \sum D_\nu]$ & $ [2.00,\ 59.4]{\rm meV}$  \\
 		\rule[14pt]{0pt}{0pt}
 		$\chi^2$ & $0.0593$ 	\\
 		\hline
 	\end{tabular}
 	\caption{Best fit of $\chi^2$ in case of NH.	 
 	}
 	\label{BP-nh}
 \end{table}
%
%
In case of {\it IH} case, we obtain the best fit value of $\chi^2$ is 0.0673 and show our input values and observables in Table~\ref{BP-ih}.
 \begin{table}[h!]
 		\centering
 	\begin{tabular}{|c|c|} \hline 
 		\rule[14pt]{0pt}{0pt}	
 		$\delta m$& $5.48$ GeV  \\ 
 		\rule[14pt]{0pt}{0pt}
		$[M_{X_1},M_{X_2},M_{X_3}]$& $[138,\ 9.24\times10^3,\ 1.63\times10^4]$ GeV  \\ 
 		\rule[14pt]{0pt}{0pt}
 		$[\tilde M_{1},\tilde M_{2}]$ &$[0.00203,\ 0.0685]$ \\
 		\rule[14pt]{0pt}{0pt}
 		$[y_{11},\ y_{22}]$ &$[0.00540,\ 0.505]$ \\
 		\rule[14pt]{0pt}{0pt}
 		$ f_L \times10^2$ & $
		\left[\begin{array}{ccc}
-0.00236 - 0.0278 i & -21.7 - 79.6 i & 0.00106 + 0.00142 i \\
6..52 - 0.529  i & -0.00651 + 0.00339  i & 0.00509 + 0.00146  i \\
22.0 + 51.7 i & 47.4 - 18.2  i & 0.0713 - 1.31 i \\
\end{array}\right]$  \\
 		\rule[14pt]{0pt}{0pt}
         	$|F|_{max}$ & $0.00171$   \\
 		\rule[14pt]{0pt}{0pt}
 	$\kappa_\nu$ & $ 1.86\times 10^{-9}$  \\
 		\rule[14pt]{0pt}{0pt}
	$[\sin\theta_{12},\ \sin\theta_{23},\ \sin\theta_{13}]$ & $ [0.554, 0.759, 0.149]$  \\
 		\rule[14pt]{0pt}{0pt}
 	$[\Delta m^2_{sol},\ \Delta m^2_{atm}] $ & $[7.51\times 10^{-5},\ 2.50\times 10^{-3}]{\rm eV^2}$  \\
 		\rule[14pt]{0pt}{0pt}
 		$[\delta_{CP},\alpha,\beta]$ & $[92.0^\circ, 21.3^\circ, 323^\circ]$ \\
 		\rule[14pt]{0pt}{0pt}
	$[m_{ee},\ \sum D_\nu]$ & $ [47.6,\ 99.2]{\rm meV}$  \\
 		\rule[14pt]{0pt}{0pt}
 		$\chi^2$ & $0.0673$ 	\\
 		\hline
 	\end{tabular}
 	\caption{Best fit of $\chi^2$ in case of IH.	 
 	}
 	\label{BP-ih}
 \end{table}
\\

\section{Summary and discussion}
We have proposed natural hierarchies among neutral fermions in a framework of inverse seesaw, imposing a $Z_3$ Tambara-Yamagami fusion rule which has recently been applied to our phenomenology. Under the symmetry, the Majorana mass terms for $N_R$ and $N_L$ are forbidden at tree level. However, their mass matrices are induced at one-loop level where the symmetry is dynamically broken. In addition, an appropriate dark matter candidate can be introduced in order to obtain such small mass terms. Finally, we have shown the best fit value of the neutrino sector for normal and inverted hierarchies referring Nufit 6.0.

\begin{acknowledgments}
 HO is supported by Zhongyuan Talent (Talent Recruitment Series) Foreign Experts Project. 
\end{acknowledgments}

\bibliography{ctma4}

\end{document}